### v1

# Sum Rate Maximization for NOMA-Assisted Uplink Pinching-Antenna Systems

Abstract
In this paper, we investigate an uplink communication scenario in which multiple users communicate with an access point (AP) employing non-orthogonal multiple access (NOMA). A pinching antenna, which can be activated at an arbitrary point along a dielectric waveguide, is deployed at the AP to dynamically reconfigure user channels. The objective is to maximize the system sum rate by jointly optimizing the pinching-antenna’s position and the users’ transmit powers. The formulated optimization problem is non-convex, and addressed using the particle swarm optimization (PSO) algorithm. For performance benchmarking, two time division multiple access (TDMA) schemes are considered: one based on the pinching antenna individually activated for each user, and the other based on the single-pinching-antenna configuration serving all users. Numerical results demonstrate that the use of the pinching antenna significantly enhances the system sum rate compared to conventional antenna architectures. Moreover, the NOMA-based scheme outperforms the TDMA-based scheme with a single pinching antenna but is outperformed by the TDMA-based approach when the pinching antenna is adaptively configured for each user. Finally, the proposed PSO-based method is shown to achieve near-optimal performance for both NOMA and TDMA with a common pinching-antenna configuration.
Index Terms:
Pinching-antenna, uplink, NOMA, and sum rate maximization.
I Introduction
Recently, flexible-antenna systems, such as fluid-antenna systems, have received significant attention due to their capability to dynamically reconfigure wireless channels [1]. By optimizing antenna positions, flexible-antenna systems have demonstrated superior performance compared to conventional fixed-location-antenna counterparts [2, 3]. However, the physical displacement of antennas in these systems is typically constrained to only a few wavelengths, which limits their ability to establish line-of-sight (LoS) links. This limitation is particularly critical in high-frequency bands, such as the millimeter-wave and terahertz spectrum, where the absence of LoS links severely degrades system performance [4, 5]. Addressing this challenge, the concept of pinching-antenna systems was introduced by NTT DOCOMO [4]. In their experimental demonstration, a plastic pinch was applied to a dielectric waveguide to enable the radiation of radio waves. By strategically positioning the pinch, strong LoS links can be established for users previously constrained to non-line-of-sight (NLoS) conditions, thereby significantly enhancing overall system performance.
In multi-user pinching-antenna systems, the signal transmitted through a given dielectric waveguide is inherently a superposition of the signals from all served users, thereby motivating the adoption of non-orthogonal multiple access (NOMA) techniques [5, 6, 7]. Specifically, the authors in [5] investigated NOMA-assisted pinching-antenna systems with the objective of maximizing the sum rate. Analytical results in [5] demonstrated that NOMA-assisted architectures outperform their orthogonal multiple access counterparts in terms of sum rate performance. However, achieving optimal system performance relies on the ability to activate pinching antennas at arbitrary positions along the waveguide — a requirement that poses practical implementation challenges. To address this, [6] proposed a low-complexity and hardware-friendly approach, where pinching-antennas were pre-installed at discrete, fixed locations prior to transmission. During operation, a subset of these antennas can be selectively activated to serve the users. In contrast to [5] and [6], which primarily focused on the sum rate maximization, the study in [7] addressed power minimization for NOMA-assisted pinching-antenna systems subject to each user’s minimum data rate requirement, and proposed an iterative power allocation algorithm. Numerical evaluations in [7] confirmed the efficiency and superior power-saving performance of pinching antennas over conventional ones.
Note that the aforementioned studies [5, 6, 7] primarily focused on downlink transmission in pinching-antenna systems. To the best of our knowledge, only a limited number of works — namely [8, 9] — explored uplink transmission. However, none of these uplink-focused studies have investigated the integration of NOMA into pinching-antenna systems. To address this gap, this paper considers a NOMA-assisted uplink scenario incorporating a single activated pinching antenna, with the objective of maximizing the system sum rate through a joint optimization of user transmit power and antenna positioning. The resulting optimization problem is reformulated into a generalized bell-shaped membership function, which is inherently non-convex. To solve it, we employ the particle swarm optimization (PSO) algorithm due to its effectiveness in handling non-convex problems. For performance benchmarking, we consider two time division multiple access (TDMA) schemes: one is to activate the pinching antenna at positions tailored to users in different time slots, and the other is to activate a single pinching antenna to serve all users. Numerical results reveal that the use of a pinching-antenna substantially improves the sum rate over traditional fixed-antenna designs. Additionally, the NOMA-based scheme outperforms the TDMA-based scheme with a common pinching-antenna, but is surpassed by the TDMA-based scheme when the antenna position is dynamically tailored to individual users.
Figure 1: Illustration of the considered uplink pinching-antenna system.
II System Model and Problem Formulation
II-A System Model
As illustrated in Fig. 1, we consider an uplink transmission scenario in which $M$ users communicate with a base station (BS) equipped with a single dielectric waveguide. We consider that a single pinching antenna is activated on the dielectric waveguide to establish LoS links with the users. Denote the user set by =$\mathcal{M}=\{1,\cdot\cdot\cdot,M\}$. To model the spatial configuration, we adopt a three-dimensional Cartesian coordinate system. Without loss of generality, the waveguide is assumed to have a length of $L$ and is positioned parallel to the $x$-axis at a fixed height $d$, as shown in Fig. 1. The users are randomly distributed within a rectangular region lying in the $x-y$ plane, with dimensions $D_{x}$ and $D_{y}$. Let $\Phi^{\text{Pin}}=(x^{\text{Pin}},0,d)$ denote the location of the pinching antenna, while $\Phi_{m}=(x_{m},y_{m},0)$ denote the coordinate of the $m$-th user, subject to $x_{m}\in[0,D_{x}]$ and $y_{m}\in[-D_{y}/2,D_{y}/2],\forall m\in\mathcal{M}$.
In this work, we consider a scenario in which all users are served simultaneously based on the principles of NOMA. At the BS, successive interference cancellation (SIC) is employed to mitigate inter-user interference. Without loss of generality, we assume that the users’ signals are decoded in an ascending order. Then, the achievable data rate of the $m$th user is given by:
$R_{m}^{\text{NOMA}}=\log_{2}\bigg{(}1+\frac{\eta P_{m}/\absolutevalue{\Phi_{m}-\Phi^{\text{Pin}}}^{2}}{\sum_{i=m+1}^{M}\eta P_{i}/\absolutevalue{\Phi_{i}-\Phi^{\text{Pin}}}^{2}+\sigma^{2}}\bigg{)},$ (1)
where $\eta=\frac{c^{2}}{16\pi^{2}f_{c}^{2}}$, with $c$ and $f_{c}$ denoting the speed of light and the carrier frequency, respectively. Additionally, $P_{m}$ represents the transmission power of the $m$th user, satisfying $P_{m}\leq P_{m}^{\max}$, with $P_{m}^{\max}$ denoting the maximum power constraint. $\sigma^{2}$ represents the power of additive white Gaussian noise at the BS.
II-B Problem Formulation
In this paper, we aim to maximize the sum rate of the aforementioned uplink NOMA system by jointly optimizing the transmit power of the users and the position of the pinching antenna. The resulting optimization problem can be formulated as follows:
$\max_{x^{\text{Pin}},P_{m}}\leavevmode\nobreak$ $\sum_{m=1}^{M}R_{m}^{\text{NOMA}}$ (2a)
s.t. $x^{\text{Pin}}\in[0,L],$ (2b)
$P_{m}\leq P_{m}^{\max},\forall m,$ (2c)
where (2b) limits the pinching antenna to the dielectric waveguide, while (2c) constrains the transmit power of each user to its maximum power.
III Proposed Solution
Problem (2) is non-convex due to the non-convex objective function (2a). To address it, we re-formulate (2a) as follows:
$\sum_{m=1}^{M}R_{m}^{\text{NOMA}}$ $=\sum_{m=1}^{M}\log_{2}\bigg{(}1+\frac{\eta P_{m}/\absolutevalue{\Phi_{m}-\Phi^{\text{Pin}}}^{2}}{\sum_{i=m+1}^{M}\eta P_{i}/\absolutevalue{\Phi_{i}-\Phi^{\text{Pin}}}^{2}+\sigma^{2}}\bigg{)}$ (3a)
$=\log_{2}\bigg{(}1+\frac{\sum_{m=1}^{M}\eta P_{m}/\absolutevalue{\Phi_{m}-\Phi^{\text{Pin}}}^{2}}{\sigma^{2}}\bigg{)},$ (3b)
where the final equality follows from the fact that the terms within the brackets in the sum-rate expression constitute a telescoping product [10].
On this basis, problem (2) can be re-written as
$\max_{x^{\text{Pin}},P_{m}}\leavevmode\nobreak$ $\log_{2}\bigg{(}1+\frac{\sum_{m=1}^{M}\eta P_{m}/\absolutevalue{\Phi_{m}-\Phi^{\text{Pin}}}^{2}}{\sigma^{2}}\bigg{)}$ (4a)
s.t. $x^{\text{Pin}}\in[0,L],$ (4b)
$P_{m}\leq P_{m}^{\max},\forall m.$ (4c)
It can be readily verified that the objective function in (4a) is a monotonically increasing function of each user’s transmit power. Therefore, to maximize the system sum rate, it is optimal for each user to transmit at its maximum power level, i.e., $P_{m}=P_{m}^{\max},\forall m\in\mathcal{M}$. With this rationale, the only remaining optimization variable is the position of the pinching-antenna, denoted by $x^{\text{Pin}}$. To facilitate further analysis, we remove the $\log(\cdot)$ function from the objective, leveraging its monotonicity, and reformulate the problem as:
$\max_{x^{\text{Pin}}}\leavevmode\nobreak$ $\sum_{m=1}^{M}P_{m}^{\max}/\absolutevalue{\Phi_{m}-\Phi^{\text{Pin}}}^{2}$ (5a)
s.t. $x^{\text{Pin}}\in[0,L].$ (5b)
It is evident that problem (5) shares the same optimal solution as the original formulation in (4), due to the monotonic nature of the logarithmic function. Furthermore, by substituting the expressions for $\Phi_{m}$ and $\Phi^{\text{Pin}}$ with its corresponding coordinates, the objective function (5a) can be re-expressed as
$\sum_{m=1}^{M}\frac{P_{m}^{\max}}{\absolutevalue{x^{\text{Pin}}-x_{m}}^{2}+y_{m}^{2}+d^{2}},$ (6)
where $x_{m}$ and $y_{m}$ are the coordinates of the $m$th user in the $x$- and $y$-axis, which are assumed known.
Each term in the objective function corresponds to a generalized bell-shaped membership function, which exhibits a symmetric bell-shaped profile [11]. As illustrated in Fig. 2, these functions are concave in the vicinity of their peak (i.e., near the center), convex in regions further from the center, and not globally convex.
Figure 2: Illustration of generalized bell-shaped membership function and its derivative.
Figure 3: One realization of the objective function with randomly generated values of $x_{m}$, $y_{m}$ and $d$, when $M=5$. The red circles denote the corresponding $x_{m}$ values.
As shown in Fig. 3, the sum of multiple symmetric bell-shaped functions may exhibit multiple local maximums, making it challenging to derive a closed-form or analytical solution. Notably, multiple local maximums can occur even within the interval between two adjacent users along the $x$-axis. This characteristic renders traditional methods, such as the bisection algorithm, unsuitable for reliably identifying the maximum between two adjacent users. Given that the problem involves a single optimization variable, a one-dimensional exhaustive search can always be employed to find the global optimum. However, the computational complexity of this approach increases with the search resolution (granularity).
As an alternative to the exhaustive search, we propose the use of PSO — a heuristic optimization technique that has demonstrated strong performance in identifying near-optimal solutions for non-convex problems [12]. The implementation of PSO in our context is as follows: We first convert the original maximization problem into a minimization problem by negating the objective function. Let $I$ denote the total number of particles in the swarm. The position of the $i$th particle at iteration $t$ is represented by $x^{\text{Pin}}_{i}(t)$, and each particle is also associated with a velocity denoted by $v^{\text{Pin}}_{i}(t)$. The particle positions are iteratively updated, where the new position of each particle at the next iteration is computed as:
$x^{\text{Pin}}_{i}(t+1)=x^{\text{Pin}}_{i}(t)+v^{\text{Pin}}_{i}(t+1).$ (7)
Simultaneously, the velocities of the particles are updated according to the following rule:
$v^{\text{Pin}}_{i}(t+1)=wv^{\text{Pin}}_{i}(t)+c_{1}r_{1}(pbest^{i}-x^{\text{Pin}}_{i}(t))+c_{2}r_{2}(gbest-x^{\text{Pin}}_{i}(t)),$ (8)
where $r_{1}$ and $r_{2}$ are random variables in the range $[0,1]$; $w\in[0,1]$ denotes the inertia weight, which controls the influence of a particle’s previous velocity; and $c_{1}$ and $c_{2}$ are known as the cognitive and social acceleration coefficients, respectively. These parameters collectively govern the balance between exploration (global search) and exploitation (local refinement). The term $pbest^{i}$ represents the personal best position discovered by the $i$th particle, while $gbest$ is the global best position found by any particle in the swarm. Both are updated simultaneously at each iteration to reflect the best solutions encountered thus far. The algorithm terminates when either the maximum number of iterations is reached, or the relative change in the global best objective function values of two consecutive iterations falls below a predefined tolerance threshold. In the context of our problem, the search space for $x^{\text{Pin}}$ is bounded. Specifically, the lower bound is set to $\min(x_{m})$, and the upper bound is given by $\min(\max(x_{m}),L)$, since the values of $x^{\text{Pin}}$ outside this interval are guaranteed to yield suboptimal results due to their distances from all user positions.
IV Two TDMA-based Benchmarks
To evaluate the performance of NOMA in pinching-antenna systems, we consider TDMA as a benchmark. In TDMA, users are served individually in distinct time slots. For simplicity and to ensure fairness, we assume equal time allocation across users, with user $m$ being served in the $m$th time slot. In this context, the position of the pinching-antenna can either be dynamically adjusted to serve each user individually or fixed to serve all users collectively, as in the NOMA case. While the former approach can potentially achieve superior performance by optimizing the antenna position for each user, it leads to more system complexity due to the need of frequent position adjustment. In the following, we analyze these two TDMA scenarios.
IV-A TDMA with the Pinching-Antenna Location Tailored to Each User
In this case, the achievable rate for user $m$ is given by
$R_{m}^{\text{TDMA}}$ $=\frac{1}{M}\log_{2}\bigg{(}1+\frac{\eta MP_{m}}{\absolutevalue{\Phi_{m}-\Phi^{\text{Pin}}_{m}}^{2}\sigma^{2}}\bigg{)}$ (9a)
$=\frac{1}{M}\log_{2}\bigg{(}1+\frac{\eta MP_{m}/\sigma^{2}}{\absolutevalue{x^{\text{Pin}}_{m}-x_{m}}^{2}+y_{m}^{2}+d^{2}}\bigg{)},$ (9b)
where $\Phi^{\text{Pin}}_{m}$ and $x^{\text{Pin}}_{m}$ denote the position of the pinching-antenna for user $m$ and its corresponding $x$-coordinate, respectively. For the purpose of a fair comparison, the factor $M$ is included in the numerator to ensure equal power budget for each user under TDMA and NOMA. Clearly, to maximize the rate, each user should operate at their maximum transmit power, i.e., $P_{m}=P_{m}^{\max}$. Moreover, $R_{m}^{\text{TDMA}}$ is maximized when |$\absolutevalue{x^{\text{Pin}}_{m}-x_{m}}^{2}$ is minimized. Given the movement constraints of the pinching-antenna, we have $x^{\text{Pin}}_{m}=\min(L,x_{m})$. That is, when $x_{m}\leq L$, $x^{\text{Pin}}_{m}=x_{m}$, meaning the pinching-antenna aligns directly with the user. Otherwise, the pinching-antenna is positioned at the closest point to the user, i.e., $x^{\text{Pin}}_{m}=L$.
IV-B TDMA with a Single Pinching-Antenna Location for All Users
In this scenario, all users share the same pinching-antenna location, as in the NOMA configuration. Consequently, the achievable rate for user $m$ can be expressed as
$R_{m}^{\text{TDMA}}$ $=\frac{1}{M}\log_{2}\bigg{(}1+\frac{\eta MP_{m}}{\absolutevalue{\Phi_{m}-\Phi^{\text{Pin}}}^{2}\sigma^{2}}\bigg{)}$ (10a)
$=\frac{1}{M}\log_{2}\bigg{(}1+\frac{\eta MP_{m}/\sigma^{2}}{\absolutevalue{x^{\text{Pin}}-x_{m}}^{2}+y_{m}^{2}+d^{2}}\bigg{)}.$ (10b)
Similar to the NOMA case, $R_{m}^{\text{TDMA}}$ exhibits a symmetric bell-shaped curve, and its summation may contain multiple local optima. Consequently, we can apply either a one-dimensional exhaustive search or PSO to identify its solution.
Figure 4: Achievable sum rate versus the maximum transmit power constraint at the users.
V Numerical Results
Numerical simulations are conducted to evaluate the performance of the proposed pinching-antenna system in comparison with the conventional fixed-antenna system, where the antenna is positioned at a fixed location of $(0,0,d)$ m. The default simulation parameters are as follows [13]: carrier frequency $f_{c}=28$ GHz, antenna height $d=3$ m, and noise power $\sigma^{2}=-90$ dBm. The number of users is set to 5, randomly deployed within a service area of $D_{x}=120$ m and $D_{y}=20$ m. The maximum transmit power for each user is constrained to $P_{m}^{\max}=25$ dBm, $\forall m$. Additionally, the length of the dielectric waveguide is set to $L=D_{x}$. All presented results are averaged over $10^{4}$ independent random realizations.
Figure 4 shows the achievable sum rate as a function of the users’ maximum transmit power. As expected, the sum rate increases for all schemes with rising transmit power. For the conventional fixed-antenna system, only the NOMA-based performance is shown, as the TDMA variant is inherently bounded by NOMA in terms of sum rate performance [14]. The results clearly demonstrate that the pinching-antenna system significantly outperforms the fixed-antenna counterpart, benefiting from its capability to dynamically reposition the antenna closer to the users. Among the pinching-antenna configurations, the NOMA-based scheme achieves a higher sum rate than the TDMA-based scheme with a single antenna location (“TDMA-Exhaustive” and “TDMA-PSO”), but is outperformed by the TDMA-based scheme when the antenna is dynamically repositioned for each user (“TDMA-Adjusted”). Furthermore, the sum rate achieved by the proposed PSO-based optimization closely matches that obtained via an exhaustive search for both NOMA and TDMA scenarios, confirming its near-optimal performance.
Figure 5 illustrates the variation of the achievable sum rate with respect to $D_{x}$, which defines the extent of the user service region along the $x$-axis. The $y$-axis range is fixed at $D_{y}=20$ m, and the dielectric waveguide length is set to $L=D_{x}$. For all considered schemes — except TDMA with user-specific pinching-antenna adjustments — the sum rate decreases as $D_{x}$ increases. This decline is attributed to greater spatial separation between users along the $x$-axis, leading to reduced channel gains. Nevertheless, the pinching-antenna systems consistently outperform the fixed-antenna baseline, with the performance gap widening as $D_{x}$ increases, thereby highlighting the advantage of antenna repositioning capabilities. Furthermore, the NOMA-based scheme continues to outperform TDMA with a common pinching-antenna, with the performance margin becoming more pronounced at larger $D_{x}$. In contrast, the TDMA scheme with adaptive pinching-antenna positioning for each user maintains a constant sum rate regardless of $D_{x}$. This is due to the ability to reposition the antenna to align with each user’s $x$-coordinate, i.e., ensuring |$\absolutevalue{x^{\text{Pin}}-x_{m}}^{2}=0$. These observations further validate the performance benefits of employing a pinching-antenna in wireless systems.
Figure 5: Achievable sum rate versus $D_{x}$, with $D_{y}=20$ m and $L=D_{x}$.
VI Conclusion
In this paper, we studied the problem of sum rate maximization in an uplink communication system employing a pinching-antenna. For the NOMA-based transmission scenario, we demonstrated that the associated optimization problem is non-convex and proposed a PSO approach to efficiently obtain a near-optimal solution. The performance of the proposed NOMA scheme was evaluated against two TDMA-based benchmarks: one based on the pinching antenna individually activated for each user, and the other based on the single-pinching-antenna configuration serving all users. Numerical simulations revealed that systems employing a pinching antenna significantly outperform conventional fixed-antenna architectures in terms of achievable sum rate. Additionally, the NOMA-based approach provides superior performance compared to TDMA with a common pinching-antenna configuration, but is outperformed by TDMA when the antenna location is adaptively optimized for each user. These findings highlight the potential of pinching-antenna systems and offer insights into their integration with multiple access techniques for enhanced spectral efficiency.
In this paper, the scenario with a single activated pinching antenna was considered. An important direction for future research is to study the performance of NOMA assisted pinching-antenna systems with multiple activated pinching antennas, where different conclusions about the comparison between NOMA and TDMA schemes might be drawn.